Compute and Compare ROE, ROA, and RNOA
Selected balance sheet and income statement information for Oracle Corporation follows. (Perform the required computations from the perspective of an Oracle shareholder.

$ millions May 31, 2015 May 31, 2014
Operating assets $62,090 $57,002
Nonoperating assets 54,368 38,819
Total assets 116,458 95,821
Operating liabilities 20,180 19,055
Nonoperating liabilities 41,958 24,097
Total liabilities 62,138 43,152
Total Oracle stockholders' equity 54,320 52,669
Total revenues 38,559
Operating income before tax 13,993
Nonoperating expense before tax 1,037
Tax expense 3,498
Net income 9,458


a. Compute return on equity (ROE)
Round answer to two decimal places (ex: 0.12345 = 12.35%)
Answer

%

b. Compute return on net assets (ROA)
Round answers to two decimal places (percentage ex: 0.12345 = 12.35%)
Answer

%

c. Compute return on net operating assets (RNOA)
Round answers to two decimal places (percentage ex: 0.12345 = 12.35%)
Answer

%
Answer #1
Ans.(a) ROE = Net Income / Average Equity
= 9,458 / (52,669+54,320)/2
= 9,458 / 53,494.5
= 17.68%
Ans.(b) ROE = Net Income / Average Assets
= 9,458 / (95,821 + 116,458)/2
= 9,458 / 106,139.5
= 8.91%
Ans.(c) RNOA = Net operating profit after tax / Average net operating assets
= 9,458 / (37,947 + 41,910)/2
= 9,458 / 39,928.5
= 23.69%
Working Note:
May 31, 2015 May 31, 2014
Operating Assets             62,090             57,002
Operating Liabilities             20,180             19,055
   NOA             41,910             37,947